# The generalized Hierarchical Gaussian Filter

### v2

@@ 5.2 Implementing the HGF’s message passing scheme
Hierarchical filtering and predictive coding are two prominent classes of hierarchical Bayesian models that cast perception as inference and model belief updates in proportion to precision-weighted prediction errors. Models from both classes are widely used, both in basic (computational) neuroscience, and for understanding mental disorders in computational psychiatry (Petzschner et al., 2017).
While the message passing architecture implied by different predictive coding models has been examined in detail and partly matched with neuroanatomy and -physiology (for overviews, see Spratling, 2019; Keller & Mrsic-Flogel, 2018; Bastos et al., 2012; Shipp, 2016), and hierarchical filtering models share many similarities with predictive coding models, it is currently not clear whether the respective inference networks would place distinct requirements on implementation (in computers or brains), or make distinct predictions about neural readouts of perceptual inference and learning. This is partly due to their non-overlapping applications: predictive coding models consider hierarchies in which higher levels affect the mean of lower levels, and they are typically used to model inference about static sensory inputs in continuous time.
@@
One noteworthy difference between the architectures is that the update equations in the HGF require a bottom-up transmission of lower-level precision estimates (Figures 3 and 4) [^4] It is not surprising that the differences between the models concern the signalling of precision: The HGF derivation explicitly includes update equations for the precision associated with beliefs - as do other hierarchical Bayesian architectures based on Markovian processes (Friston et al., 2013). In contrast, most predictive coding schemes only focus on the optimization of the first moment (mode or expectation) of the posterior distribution for perceptual inference (Friston, 2005; Bogacz, 2017), although the variational approach does allow approximation of the full posterior distribution including its variance.. This is interesting, given that recent neuroanatomical studies point to additional pathways besides the classical forward (from lower-level supragranular to higher-level granular layers) and backward (from higher-level infragranular to lower-level extragranular layers) connections (Markov et al., 2013, 2014). Our architecture is compatible with an ascending connection within supragranular layers (for bottom-up communication of lower-level precision) that runs in parallel to a descending connection within these layers (for top-down modulation of PEs by higher-level precision), reminiscent of the “cortical counter streams” identified by these studies. We hope to capitalize on methodological advancements in high-resolution laminar fMRI (Stephan et al., 2019; Haarsma et al., 2022) in future studies to test these predictions.
@@ 5.3 Dynamics within versus across time steps
Although potential neurobiological implementations of approximate Bayesian inference in the brain are still hotly debated (Knill & Pouget, 2004; Aitchison & Lengyel, 2017), a growing body of literature suggests that predictive coding-like architectures can account for a large range of neurophysiological findings in perception research (for a recent overview, see Walsh et al., 2020), making these architectures particularly relevant for understanding human perception and inference. It would thus be interesting to examine in detail the commonalities and potential differences between message-passing schemes implied by different forms of coupling in HGF models and classical predictive coding schemes, as we have started to do here.
@@
Importantly, however, direct comparisons of the two models are further complicated by differences in each model’s concept of time: while the HGF captures belief updates across time steps in discrete time (i.e., across sequentially arriving sensory inputs), predictive coding describes the evolution of beliefs in continuous time and, typically, in response to static sensory input (Rao & Ballard, 1999; Friston, 2005) [^5] In fact, the standard form of the generative model in predictive coding does not contain temporal dynamics of the hidden states, but see Friston (2008, 2010) for extensions.. Differential equations capture the evolution of beliefs and predictions errors and are used to simulate perceptual inference, starting with new input to the lowest level of the belief hierarchy, and ending when the ensuing PEs have been reconciled, i.e., a stable new posterior belief has emerged (Bogacz, 2017). This can be used to make predictions about neural activity that can be compared against measurements. While, to our knowledge, existing predictive coding models have not been fit to data, the simulated neural dynamics display many features that are observed in real data, such as oscillatory tendencies, even in very small networks (Bogacz, 2017). We refer to these simulations as within-step dynamics of belief updating.
@@
On the other hand, the generative model of the HGF represents a Markovian process in discrete time; in the inference model, one-step update equations, derived based on a mean field approximation to the full Bayesian solution, quantify the change from prior to posterior on all levels of the belief hierarchy. The model is thus examined in sequential input settings to capture step-by-step learning - in other words, across-step dynamics of belief updating. The model provides an approximately Bayes-optimal solution to stepwise belief updating, useful for ideal observer analyses (Stefanics et al., 2018; Weber et al., 2020, 2022; Hauke et al., 2022), but the parameters of the HGF can also be fit to behavioural responses of individual participants. Model-derived agent-specific trajectories of predictions and PEs have proven particularly useful for identifying potential neural and physiological correlates of computational quantities in empirical data (Iglesias et al., 2013b; Vossel et al., 2015; de Berker et al., 2016; Diaconescu et al., 2017; Weilnhammer et al., 2018; Katthagen et al., 2018; Palmer et al., 2019; Deserno et al., 2020; Henco et al., 2020; Cole et al., 2020; Lawson et al., 2021; Hein et al., 2021; Hein & Herrojo Ruiz, 2022; Harris et al., 2022; Fromm et al., 2023). In the future, we propose explicitly to consider potential within-step dynamics of belief updates compatible with the update equations of the HGF (see Weber, 2020, for a first attempt). Such equations could for example be derived by treating the HGF posterior values of all nodes (quantities) as the equilibrium point towards which all dynamics must converge (inspired by Bogacz, 2017). Establishing equations for within-step belief updating dynamics under the HGF might allow for empirical tests of the proposed architecture in a two-step procedure: first, individual trajectories of predictions and prediction errors are inferred from observed behaviour by fitting the HGF to participants’ responses, second, these stepwise point estimates are subsequently used to simulate expected continuous-time neuronal responses according to the differential equations.
@@
A modular implementation of the generalized HGF makes it easy to build large networks with considerable hierarchical depth, opening up exciting possibilities of applying this model architecture in machine learning applications and comparing its performance to alternative neural network architectures (Whittington & Bogacz, 2017; Millidge et al., 2022; Song et al., 2024). When applied as a model of human cognition, however, an increase in model complexity must be matched by sufficiently rich data to successfully fit the model. In the generalised HGF, a branching out into the higher levels of the belief hierarchy (multiple parent nodes) will typically require some amount of branching out in the other direction (multiple child nodes) to disambiguate beliefs about hidden states further up. Moreover, in this paper, we use the original update equations for volatility parents from Mathys et al. (2011). These are derived from a particular quadratic approximation to the variational energy of the parent node which can sometimes lead to a (logically impossible) negative value for the posterior precision of that node. In most cases, this can be circumvented by choosing appropriate priors. However, it can also be avoided entirely by adapting the details of the quadratic approximation. Such adaptations will be included in the generalized HGF as soon as they are published separately.
@@
In summary, we have presented a generalization of the HGF that extends its scope of hierarchical inference mechanisms to include cross-level couplings as proposed by predictive coding. Furthermore, we have demonstrated how this extension can be cast as a modular architecture that allows for flexible changes to a model without having to re-derive update equations. We hope that the availability of these developments as open source software will expand the toolkit of computational psychiatry and facilitate future investigations of perceptual inference in health and disease.

### v3

@@ 5.2 Implementing the HGF’s message passing scheme
Hierarchical filtering and predictive coding are two prominent classes of hierarchical Bayesian models that cast perception as inference and model belief updates in proportion to precision-weighted prediction errors. Models from both classes are widely used, both in basic (computational) neuroscience, and for understanding mental disorders in computational psychiatry (petzschner2017).
While the message passing architecture implied by different predictive coding models has been examined in detail and partly matched with neuroanatomy and -physiology (for overviews, see spratling2019; keller2018; bastos2012; shipp2016), and hierarchical filtering models share many similarities with predictive coding models, it is currently not clear whether the respective inference networks would place distinct requirements on implementation (in computers or brains), or make distinct predictions about neural readouts of perceptual inference and learning. This is partly due to their non-overlapping applications: predictive coding models consider hierarchies in which higher levels affect the mean of lower levels, and they are typically used to model inference about static sensory inputs in continuous time (for a recent extension to dynamic inputs, see millidge2024).
@@
One noteworthy difference between the architectures is that the update equations in the HGF require a bottom-up transmission of lower-level precision estimates (Figures 3 and 4) [^4] It is not surprising that the differences between the models concern the signalling of precision: The HGF derivation explicitly includes update equations for the precision associated with beliefs - as do other hierarchical Bayesian architectures based on Markovian processes (Friston2013). In contrast, most predictive coding schemes only focus on the optimization of the first moment (mode or expectation) of the posterior distribution for perceptual inference (Friston2005; bogacz2017), although the variational approach does allow approximation of the full posterior distribution including its variance.. This is interesting, given that recent neuroanatomical studies point to additional pathways besides the classical forward (from lower-level supragranular to higher-level granular layers) and backward (from higher-level infragranular to lower-level extragranular layers) connections (Markov2013; Markov2014). Our architecture is compatible with an ascending connection within supragranular layers (for bottom-up communication of lower-level precision) that runs in parallel to a descending connection within these layers (for top-down modulation of PEs by higher-level precision), reminiscent of the “cortical counter streams” identified by these studies. We hope to capitalize on methodological advancements in high-resolution laminar fMRI (Stephan2019; haarsma2022) in future studies to test these predictions.
@@ 5.3 Dynamics within versus across time steps
Although potential neurobiological implementations of approximate Bayesian inference in the brain are still hotly debated (knill2004; aitchison2017), a growing body of literature suggests that predictive coding-like architectures can account for a large range of neurophysiological findings in perception research (for a recent overview, see Walsh2020), making these architectures particularly relevant for understanding human perception and inference. It would thus be interesting to examine in detail the commonalities and potential differences between message-passing schemes implied by different forms of coupling in HGF models and classical predictive coding schemes, as we have started to do here.
@@
Importantly, however, direct comparisons of the two models are further complicated by differences in each model’s concept of time: while the HGF captures belief updates across time steps in discrete time (i.e., across sequentially arriving sensory inputs), predictive coding describes the evolution of beliefs in continuous time and, typically, in response to static sensory input (Rao1999; Friston2005) [^5] In fact, the standard form of the generative model in predictive coding does not contain temporal dynamics of the hidden states, but see friston2008; Friston2010 for extensions.. Differential equations capture the evolution of beliefs and predictions errors and are used to simulate perceptual inference, starting with new input to the lowest level of the belief hierarchy, and ending when the ensuing PEs have been reconciled, i.e., a stable new posterior belief has emerged (bogacz2017). This can be used to make predictions about neural activity that can be compared against measurements. While, to our knowledge, existing predictive coding models have not been fit to data, the simulated neural dynamics display many features that are observed in real data, such as oscillatory tendencies, even in very small networks (bogacz2017). We refer to these simulations as within-step dynamics of belief updating. Furthermore, predictive coding networks have recently been rediscovered as potent candidates for solving machine learning tasks (e.g., the classification or generation of static images, sun2020; song2020) and even a potential alternative to the backpropagation algorithm (song2024; millidge2022predictive; whittington2017).
@@
On the other hand, the generative model of the HGF represents a Markovian process in discrete time; in the inference model, one-step update equations, derived based on a mean field approximation to the full Bayesian solution, quantify the change from prior to posterior on all levels of the belief hierarchy. The model is thus examined in sequential input settings to capture step-by-step learning - in other words, across-step dynamics of belief updating. The model provides an approximately Bayes-optimal solution to stepwise belief updating, useful for ideal observer analyses (stefanics_visual_2018; weber2020; weber2022; hauke_aberrant_2022), but the parameters of the HGF can also be fit to behavioural responses of individual participants. Model-derived agent-specific trajectories of predictions and PEs have proven particularly useful for identifying potential neural and physiological correlates of computational quantities in empirical data (iglesias2013; vossel_cortical_2015; de_berker_computations_2016; diaconescu2017social; weilnhammer_neural_2018; katthagen_modeling_2018; palmer_sensorimotor_2019; deserno_volatility_2020; henco_bayesian_2020; cole_atypical_2020; lawson_computational_2021; hein_state_2021; hein_state_2022; harris_task-evoked_2022; fromm_belief_2023). In the future, we propose to explicitly consider potential within-step dynamics of belief updates compatible with the update equations of the HGF (see weber2020diss, for a first attempt). Such equations could for example be derived by treating the HGF posterior values of all nodes (quantities) as the equilibrium point towards which all dynamics must converge (inspired by bogacz2017). Establishing equations for within-step belief updating dynamics under the HGF might allow for empirical tests of the proposed architecture in a two-step procedure: first, individual trajectories of predictions and prediction errors are inferred from observed behaviour by fitting the HGF to participants’ responses, second, these stepwise point estimates are subsequently used to simulate expected continuous-time neuronal responses according to the differential equations.
@@
A modular implementation of the generalized HGF makes it easy to build large networks with considerable hierarchical depth, opening up exciting possibilities of applying this model architecture in machine learning applications and comparing its performance to alternative neural network architectures (whittington2017; millidge2022; song2024). When applied as a model of human cognition, however, an increase in model complexity must be matched by sufficiently rich data to successfully fit the model. In the generalised HGF, a branching out into the higher levels of the belief hierarchy (multiple parent nodes) will typically require some amount of branching out in the other direction (multiple child nodes) to disambiguate beliefs about hidden states further up.
@@
In summary, we have presented a generalization of the HGF that extends its scope of hierarchical inference mechanisms to include cross-level couplings as proposed by predictive coding. The result is a new class of artificial neural networks that incorporate computational principles of two popular hierarchical Bayesian inference schemes: HGF and predictive coding. Furthermore, we have demonstrated how this extension can be cast as a modular architecture that allows for flexible changes to a model without having to re-derive update equations, thus allowing for a modular construction of complex hierarchically structured ANNs. Finally, the availability of these developments as open source software expands the toolkit of computational psychiatry and we hope that it will facilitate future investigations of perceptual inference in health and disease.